OpenStudy (anonymous):

how does sin(-13pi/4) = -5pi/4 we were asked to find the exact value of sin(-13pi/4) how does sin(-13pi/4) = -5pi/4 we were asked to find the exact value of sin(-13pi/4) @Mathematics

OpenStudy (jamesj):

sin(-13pi/4) = sin(-5pi/4) = sin (3pi/4) because sin(x + 2pi) = sin(x)

OpenStudy (anonymous):

could you demonstrate that pls im having a hard time understanding

OpenStudy (anonymous):

-13pi/4 is over one revolution so add 2pi to it.

OpenStudy (jamesj):

Think about the unit circle. Go around one more complete revolution then you're still at the same point on the unit circle, but the angle has changed by 2pi

OpenStudy (anonymous):

\[-\frac{13 \pi}{4} + \frac{8\pi}{4} = -\frac{5\pi}{4}\]

OpenStudy (anonymous):

oh tyler...now i see

OpenStudy (anonymous):

and also could you demonstrate how cos(5pi/3)= -sqrt2/2 and cos(-270) = 0

myininaya (myininaya):

\[\sin(\frac{-13 \pi}{4})=\sin(\frac{-4}{4}\pi+\frac{-9}{4}\pi)=\sin(-\pi-\frac{4}{4}\pi-\frac{5}{4} \pi)=\sin(-\pi-\pi-\frac{5}{4} \pi)\] \[\sin(-2\pi-\frac{5}{4}\pi)=\sin(-(2\pi+\frac{5}{4}\pi))\]

OpenStudy (anonymous):

cos(5pi/3) should be 1/2 not what you said

OpenStudy (anonymous):

and change -270 into positive degrees so you do 360 + -270 = 90 and cos(90) = 0

OpenStudy (anonymous):

oh thanks for the help again Tyler. how have you been?

OpenStudy (anonymous):

I've been good.. how about you?

OpenStudy (anonymous):

i have been good as well...preparing for the test. its on wednesday. I almost have everything down. I may have one or two questions that I need help, will you be around this evening?

OpenStudy (anonymous):

yeah i can help if you want

OpenStudy (anonymous):

ok i need help with inverses. first I have tan(x/2) + (pi/6) and i need to find the inverse

OpenStudy (jamesj):

Note by the way that even if we have shown sin(-13pi/4) = sin (3pi/4), that's not the final answer. You need to evaluate sin(3pi/4)

OpenStudy (anonymous):

oh ok thanks for letting me know

OpenStudy (anonymous):

Ok inverses.. just a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

to be honest, ive never had to do that before so im not sure on how to do it.

OpenStudy (anonymous):

oh ok thats fine.

OpenStudy (jamesj):

cos x = 1 - 2 sin^2 (x/2) hence \[ \sin(x/2) = \pm \sqrt{1 - \cos x} \]

OpenStudy (anonymous):

remember last time we did amp, domain range,period,etc...? there was one thing else i need to ask...how do you find horizontal shift

OpenStudy (anonymous):

horizontal shift is the phase shift

OpenStudy (anonymous):

if you're given y=3csc(x + pi/4)+ 2

OpenStudy (jamesj):

3pi/4 is in the first quadrant, so we will want sin(3pi/4) to be positive. Thus: \[ \sin(3\pi/4) = \sqrt{1 - cos(3\pi/2)} \]

OpenStudy (anonymous):

James do you know how to do this?? first I have tan(x/2) + (pi/6) and i need to find the inverse

OpenStudy (anonymous):

pi/4 is the horizontal shift aka phase shift

OpenStudy (anonymous):

thanks my teacher got the horizontal shift as -pi/2 for y=-2cos (2x + pi) - 1...im guessing she made a mistake right?

OpenStudy (anonymous):

I would think.. because horizontal means going left and right.. so the horizontal shift would be -pi

OpenStudy (anonymous):

yeah that makes sense.

OpenStudy (anonymous):

Horizontal Shifts To shift a graph horizontally, a constant must be added to the function within parentheses--that is, the constant must be added to the angle, not the whole function.

OpenStudy (anonymous):

so james do you know how to find that inverse?

OpenStudy (anonymous):

so yeah the teacher made a mistake

OpenStudy (anonymous):

well hold on.. i think your teacher factored out a 2

OpenStudy (anonymous):

\[y = -2\cos2(x + \frac{\pi}{2})\]

OpenStudy (anonymous):

hmm thats what it now looks like

OpenStudy (anonymous):

so how did she get 2 as the denominator in the phase shift

OpenStudy (anonymous):

factoring something out is like dividing by that number so if you have \[(2x + \pi)\] and you factor out a 2 you do \[(\frac{2x}{2} + \frac{\pi}{2})\] and simplify \[(x + \frac{\pi}{2})\]

OpenStudy (anonymous):

oh now i see

OpenStudy (jamesj):

sorry, what are you asking now?

OpenStudy (jamesj):

if you've got it I'll bugger off

OpenStudy (anonymous):

first I have tan(x/2) + (pi/6) and i need to find the inverse

OpenStudy (anonymous):

james this is what we needed help with first I have tan(x/2) + (pi/6) and i need to find the inverse

OpenStudy (jamesj):

So if y = tan(x/2) + pi/6, what is x?

OpenStudy (anonymous):

no, what is the inverse of y=tan(x/2) + pi/6

OpenStudy (jamesj):

We just say the same thing. Assuming that \( -\pi/2 < x < \pi/2 \), then \[ x = 2 \arctan(y - \pi/6) \]

OpenStudy (jamesj):

Ugh, I can't type today: "We just said the same thing."

OpenStudy (jamesj):

And actually we need x/2 to be in that range -pi/2 to pi/2, hence -pi < x < pi

OpenStudy (anonymous):

ok james i have two more what about f(x) arcos(x/2)-7 f(x)= sin(3x)

OpenStudy (anonymous):

thanks for the help btw

OpenStudy (jamesj):

Let y = arccos(x/2) - 7. Now solve for x. That gives you the inverse function. The only wrinkle is to think carefully about the domain and range.

OpenStudy (anonymous):

2cos(x+7)??

OpenStudy (anonymous):

oh ok thanks for the help to both of your...can i get a help with the last one f(x)= sin(3x)

OpenStudy (anonymous):

\[\frac{1}{3}\arcsin(x)\]

OpenStudy (anonymous):

thanks tyler

OpenStudy (anonymous):

yeah lol.. im getting the hang of these lol

OpenStudy (anonymous):

lol...now i just have basic questions. if intervals = period/4 and my period is 2pi/3...how do i end up with pi/6

OpenStudy (anonymous):

Im confused :/ lol

OpenStudy (anonymous):

oh wait no im not.

OpenStudy (anonymous):

you do \[\frac{\frac{2\pi}{3}}{4}\]

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

\[\frac{2\pi}{3} * \frac{1}{4}\] which equals \[\frac{2\pi}{12}\] which equals \[\frac{\pi}{6}\]

OpenStudy (anonymous):

and how do you get 1/2 of 2pi/3 to equal pi/3

OpenStudy (anonymous):

\[\frac{\frac{2\pi}{3}}{2}\]

OpenStudy (anonymous):

\[\frac{2\pi}{3} * \frac{1}{2}\]\[\frac{2\pi}{6}\] \[\frac{\pi}{3}\]

OpenStudy (anonymous):

woow man thanks for the help i greatly appreciate it.

OpenStudy (anonymous):

no problem. Do you understand??

OpenStudy (anonymous):

yes very well.

OpenStudy (anonymous):

cool :D

OpenStudy (anonymous):

Anything else??

OpenStudy (anonymous):

thats it. maybe if i come across something later on that i dont understand as im studying. but those were the only ones i couldnt comprehend

OpenStudy (anonymous):

Cool, good luck on your test :)

OpenStudy (anonymous):

thanks for the help once again

OpenStudy (anonymous):

No problem. See you around sometime!! :D

OpenStudy (anonymous):

you too